OpenStudy (anonymous):

Divide. Give the quotient and remainder. 45÷8

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

how many times does 8 go in to 45. Then the rest is the remainder

OpenStudy (anonymous):

Whats the highest multiple of 8 that is below 45 and then subtract the "remainder" or difference

OpenStudy (anonymous):

so the quotient is 5 and remainder is 5?

jimthompson5910 (jim_thompson5910):

You are correct 45/8 = 5 remainder 5 8 goes into 45 five whole times (8*5 = 40) the amount left over is 45-40 = 5 which is your remainder

OpenStudy (skullpatrol):